The combination of two bar magnets makes 10 oscillations per second in an oscillation magnetometer when like pole are tied together and 2 oscillations per second when unlike poles are tied together. Find the ratio of the magnetic of the magnets. Neglect any induced magnetism.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. B base H = 24 × 10^–6 T T base 1 = 0.1’ B = B base H – B base wire = 2.4 × 10^–6 – μ base 0/2π i/r = 24 * 10^-6 – 2 * 10^-7 *18/0.2 = (24 - 10) * 10^-6 = 14 * 10^-6 T = 2π√I/MB base H T base 1/T base 2 = √B/B base H ⇒ 0.1/T base 2 = √14 * 10^-6/24 * 10^-6 ⇒ (0.1/T base 2)^2 = 14/24 ⇒ 0.01* 14/24 ⇒ T base 2 = 0.076